Two identical thin rods of length L (and equal mass) are joined at right angles to form an L-shaped object. This object is balanced in a vertical plane on top of a sharp edge. The L-shaped object oscillates as a physical pendulum when it is slightly deflected (small angle). Starting with T=2pi(I/mgd)^(1/2), derive an expression for the frequency of the oscillation. Write the result in terms of thequantities given in the problem and, perhaps, constants (e.g., π, ⅓, g, ...).
1
Expert's answer
2020-02-10T09:24:15-0500

The mass of the rod = m

Length of the one rod = L

As per the question, if it is joining at "90^\circ"

Then, Moment of the inertia of the combined system = "\dfrac{mL^2 }{3}+\dfrac{mL^2}{3}=\dfrac{2mL^2}{3}"

The effective length of the pendulum ="L\cos45^\circ=\dfrac{L}{\sqrt{2}}"

Now, we know that time period "T=2\pi\sqrt{\dfrac{I}{mgl}}"

"T=2\pi\sqrt{\dfrac{2mL^2\sqrt{2}}{3\times 2mgL}}"

"\Rightarrow T=2\pi\sqrt{\dfrac{L\sqrt{2}}{3\times g}}"

"\Rightarrow T=2 \pi\sqrt{\dfrac{\sqrt{2}L}{3g}}"